OpenStudy (anonymous):

find x intertcepts of y=x^2+2x-8

OpenStudy (anonymous):

x^2+2x-8 = (x+4) (x-2) x = -4, 2

OpenStudy (anonymous):

y=0 is that my first step?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

then factor and find the x values as moneybird did

OpenStudy (anonymous):

(x-2)(x+4) -2,4?

OpenStudy (anonymous):

2,-4?

OpenStudy (anonymous):

2 ,-4 yes

OpenStudy (anonymous):

does it matter what way we write it?

OpenStudy (anonymous):

no you can write it as -4, 2 or wtite 'x=2 or -4' - means same thing

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

i have soo much to check ugh